# Enabling Automatic Differentiation with Mollified Graph Neural Operators

### v1

@@ Appendix E Choices of Hyperparameters
where $\varepsilon$ is a small positive number for numerical stability, and the constant $C$ is a scaling constant taking into account the measure an dimensions of the data, to ensure that the loss is averaged correctly across the spatial dimensions.
E.1 Burgers’ Equation
For Burger’s equation, the input initial condition $u_{0}(x)$ given on a regular spatial grid in the domain $[0,1]^{2}$ is first duplicated along the temporal dimension to obtain a 2D regular grid of resolution $128\times 26$, and then passed through a 2D FNO and a mollified GNO to produce a predicted function $v=(\mathcal{G}_{m\mathrm{GNO}}\ \circ\ \mathcal{G}_{\mathrm{FNO}})(u_{0})$ approximating the solution $u(x,t)$.
For this experiment, the 2D FNO has $4$ layers, each with $26$ hidden channels and $(24,24)$ Fourier modes, and we used a a Tucker factorization with rank $0.6$ of the weights. The $m$GNO uses the half_cos weighting function with a radius of $0.1$, and a $2$-layer MLP with $[64,64]$ nodes. The resulting model has $1,019,569$ trainable parameters, and was trained in PyTorch for $10,\!000$ epochs using the Adam optimizer with learning rate $0.002$ and weight decay $10^{-6}$, and the ReduceLROnPlateau scheduler with factor $0.9$ and patience $50$.
E.2 Nonlinear Poisson Equation
For the nonlinear Poisson equation, the mesh coordinates within $[-1.4,1.4]^{2}$, signed distance functions, source terms, and boundary conditions, are passed through $\mathcal{G}_{m\mathrm{GNO}}^{decoder}\circ\mathcal{G}_{\mathrm{FNO}}\circ\mathcal{G}_{\mathrm{GNO}}^{encoder}$ model to produce an approximation to the solution $u$.
For this experiment, the input GNO has a radius of $0.16$, and a $3$-layer MLP with $[256,512,256]$ nodes. The 2D FNO has $4$ layers, each with $64$ hidden channels and $(20,20)$ Fourier modes, and acts on a latent space of resolution $64\times 64$. The output $m$GNO uses the half_cos weighting function with a radius of $0.175$, and a $3$-layer MLP with $[512,1024,512]$ nodes. The resulting model has $8,691,972$ trainable parameters, and was trained in PyTorch for $300$ epochs using the Adam optimizer with learning rate $0.0001$ and weight decay $10^{-6}$, and the ReduceLROnPlateau scheduler with factor $0.9$ and patience $2$. We used $7000$ samples for training and $3000$ samples for testing.
E.3 Hyperelasticity Equation
For the hyperelasticity equation, the mesh coordinates within $[0,1]^{2}$ and signed distance functions are passed through a $\mathcal{G}_{m\mathrm{GNO}}^{decoder}\circ\mathcal{G}_{\mathrm{FNO}}\circ\mathcal{G}_{\mathrm{GNO}}^{encoder}$ model to produce an approximation to the solution $u$ to the hyperelasticity equations.
For this experiment, the input GNO has a radius of $0.05625$, and a $3$-layer MLP with $[128,256,128]$ nodes. The 2D FNO has $4$ layers, each with $64$ hidden channels and $(20,20)$ Fourier modes, and acts on a latent space of resolution $32\times 32$. The output $m$GNO uses the half_cos weighting function with a radius of $0.1125$, and a $3$-layer MLP with $[1024,2048,1024]$ nodes. The resulting model has $11,678,211$ trainable parameters, and was trained in PyTorch for $400$ epochs using the Adam optimizer with learning rate $0.0001$ and weight decay $10^{-6}$, and the ReduceLROnPlateau scheduler with factor $0.9$ and patience $10$. We used $1000$ samples for training and $1000$ samples for testing.
E.4 Airfoil Inverse Design
For the airfoil design forward problem, the $220\times 50$ mesh point locations within $[-40,40]^{2}$ and signed distance functions are passed through a differentiable mollified GINO model of the form $\mathcal{G}_{m\mathrm{GINO}}=\mathcal{G}_{m\mathrm{GNO}}^{decoder}\circ\mathcal{G}_{\mathrm{FNO}}\circ\mathcal{G}_{m\mathrm{GNO}}^{encoder}$ to produce an approximation of the pressure field $p$.
For this experiment, the input $m$GNO uses the half_cos weighting function with a radius of $2$, and a $3$-layer MLP with $[128,128,128]$ nodes. The 2D FNO has $3$ layers, each with $16$ hidden channels and $(36,36)$ Fourier modes, and acts on a latent space of resolution $64\times 64$. The output $m$GNO uses the half_cos weighting function with a radius of $6$, and a $3$-layer MLP with $[128,128,128]$ nodes. The resulting model has $1,162,546$ trainable parameters, and was trained in PyTorch for $750$ epochs using the Adam optimizer with learning rate $0.0001$ and weight decay $10^{-9}$, and the ReduceLROnPlateau scheduler with factor $9$ and patience $5$. We used $1000$ samples for training and $200$ samples for testing.
Appendix F Effect of Using Different Training Losses
(a)(b)(c)
Figure 15: (a) Prediction of a $m$GINO model trained using data loss only for the nonlinear Poisson equation. This is the predicted solution evaluated at a high resolution on a small patch centered at a latent query point of the $m$GNO. We see that the prediction exhibits discontinuities that coincide with the circles of radius $r$ (blue lines) centered at the neighboring latent query points. (b)(c) Norm of the gradient of the predicted solution shown in (a), computed using automatic differentiation (in (b)) and finite differences (in (c)).
Figure 16: Comparison between the prediction of a $m$GINO model (left) trained using a hybrid loss, $\mathcal{L}=\mathcal{L}_{\text{data}}+\lambda\mathcal{L}_{\text{physics}}$ and the corresponding ground truth solution (right) for the nonlinear Poisson equation. The prediction and ground truth solutions are evaluated at a high resolution on a small patch centered at a latent query point of the $m$GNO. Using the physics loss regularized the higher resolution prediction, and removed the visible discontinuity patterns shown in Figure 15(a).
Appendix G Finite Differences on Point Clouds
On regular grid, standard well-known stencils formulas are available to compute first-order and higher-order derivatives using finite differences. However, on arbitrary point clouds, the varying distances between points have to be taken into account and a different stencil is needed for each point at which the derivatives need to be computed. We detail below one strategy to obtain these stencil formulas in 2D.
We consider the case where we have an arbitrary point 2D cloud with $N$ points, $\{(x_{i},y_{i})\}_{i=1}^{N}$, and suppose that the function values $\{f(x_{i},y_{i})\}_{i=1}^{N}$ of a function $f$ are known at these points. The goal is to approximate partial derivatives of $f$ at any other point $(\tilde{x},\tilde{y})$ in the domain. We start with first-order derivatives, and write them as a finite difference with unknown stencil coefficients:
@@ Appendix G Finite Differences on Point Clouds
To find the stencil coefficients $c_{i}^{(x)}$ and $c_{i}^{(y)}$, we enforce that the approximation holds exactly for the functions $1$, $x$ and $y$, that is, we enforce that the approximation holds true for any polynomial of degree 1 in 2D. This results in the following systems of equations for the stencil coefficients $c_{i}^{(x)}$ and $c_{i}^{(y)}$,
$\sum_{i=1}^{N}c_{i}^{(x)}=0,\qquad\sum_{i=1}^{N}c_{i}^{(x)}(x_{i}-\tilde{x})=1,\qquad\sum_{i=1}^{N}c_{i}^{(x)}(y_{i}-\tilde{y})=0,$
@@
$\sum_{i=1}^{N}c_{i}^{(y)}=0,\qquad\sum_{i=1}^{N}c_{i}^{(y)}(x_{i}-\tilde{x})=0,\qquad\sum_{i=1}^{N}c_{i}^{(y)}(y_{i}-\tilde{y})=1.$
@@
$A\mathbf{c}^{(x)}=\mathbf{b}^{(x)},\qquad\qquad A\mathbf{c}^{(y)}=\mathbf{b}^{(y)},$
@@
$A=\begin{bmatrix}1&1&\cdots&1\\ x_{1}-\tilde{x}&x_{2}-\tilde{x}&\cdots&x_{N}-\tilde{x}\\ y_{1}-\tilde{y}&y_{2}-\tilde{y}&\cdots&y_{N}-\tilde{y}\end{bmatrix},$
@@
$\mathbf{c}^{(x)}=\begin{bmatrix}c_{1}^{(x)}\\ ...\\ c_{N}^{(x)}\end{bmatrix},\qquad\quad\mathbf{c}^{(y)}=\begin{bmatrix}c_{1}^{(y)}\\ ...\\ c_{N}^{(y)}\end{bmatrix},\qquad\quad\mathbf{b}^{(x)}=\begin{bmatrix}0\\ 1\\ 0\end{bmatrix},\qquad\quad\mathbf{b}^{(y)}=\begin{bmatrix}0\\ 0\\ 1\end{bmatrix}.$
@@
$\mathbf{c}^{(x)}=(A^{\top}\!A)^{-1}A^{\top}\mathbf{b}^{(x)},\qquad\qquad\mathbf{c}^{(y)}=(A^{\top}\!A)^{-1}A^{\top}\mathbf{b}^{(y)}.$
Plugging these coefficients in Equation 27 gives the desired approximation to the first-order derivatives at $f$.
We emphasize that this procedure needs to be repeated for every point $(\tilde{x},\tilde{y})$ at which the derivatives need to be evaluated since the location of the point affects the entries of the matrix $A$. Note that it is not necessary and not recommended to use all $N$ points to approximate the derivatives, and one should instead identify a subset of nearest neighbors from which the finite differences can be computed.
@@
Appendix H Training times of 2D $m$GNOs with Autograd and Finite Differences (FD)
@@
We compare the training times per epoch when training 2D $m$GNOs with $\mathcal{L}_{\mathrm{Poisson}}$ using autograd, finite differences on uniform grids, and finite differences on non-uniform (NU) grids. Figures 17 and 18 display the training times per epoch versus the latent resolution and output resolution, respectively.
@@
As expected, we see from Figure 17 that autograd is more expensive than FD on a uniform grid at fixed latent and output resolutions. We can also see, by looking at individual columns corresponding to fixed latent space resolutions (i.e. the same model architecture), how FD at higher output resolutions compare to autograd at lower output resolutions. In particular, non-uniform finite differences are often more expensive than autograd at a slightly lower resolution. Figure 18 quantifies how running time increases when the latent space resolution of the model increases.
@@
Figure 17: Time per epoch when training 2D $m$GNOs with $\mathcal{L}_{\mathrm{Poisson}}$ using autograd, finite differences on uniform grids, and finite differences on non-uniform (NU) grids. We display the training times per epoch at different latent space resolutions (x-axis) and different output resolutions (the numbers next to the points denote the numbers of points along each dimension).
@@
Figure 18: Time per epoch when training 2D $m$GNOs with $\mathcal{L}_{\mathrm{Poisson}}$ using autograd, finite differences on uniform grids, and finite differences on non-uniform (NU) grids. We display the training times per epoch at different output resolutions (x-axis) and different latent space resolutions (the numbers at the end of the curves denote the numbers of points along each dimension).
Appendix I Ablation study on GNO radius and weighting function
We investigate how performance changes as the GNO radius changes for GINO models, and also compare the performance of the different weighting functions (9)-(12) on the nonlinear Poisson equation. Note that the radius cutoff used in the weighting functions of $m$GNOs is the same as the GNO radius to avoid the additional cost of carrying a second neighbor search. The results are displayed in Table 5.
The GNO radius $r$ is an important hyperparameter to tune. It needs to be large enough so that the ball $\text{B}_{r}(x)$ contains sufficiently many other latent query points, and a value too large will lead to prohibitive computational and memory costs. We denote the number of latent query points in $\text{B}_{r}(x)$ by #$\#|\text{B}_{r}(x)|$. For the nonlinear Poisson equation, we are using a regular 2D latent space grid, so $\text{B}_{r}(x)$ will only contain a single latent query point ($x$ itself) until $r$ is large enough for $\text{B}_{r}(x)$ to contain one extra latent query point in each direction. $\text{B}_{r}(x)$ will progressively contain more latent query points by thresholds as the radius increases further together with computational time and memory cost. Table 5 shows the results obtained with different weighting functions for values of $r$ such that #$\#|\text{B}_{r}(x)|=3^{2},5^{2},7^{2},9^{2},15^{2}$. When the radius is too small, performance is poor. On the other hand, when $r$ becomes too large, the computational and memory costs significantly increase while performance deteriorates. The best performance was achieved with a radius for which #$\#|\text{B}_{r}(x)|=7^{2}$, and $w_{\text{half\_cos}}$ and $w_{\text{quartic}}$ led to the best performance.
Table 5: Validation MSE of $m$GINO models, trained on only PDE loss, with different GNO radius values and weighting functions.
bump half_cos quartic octic
#$\#|\text{B}_{r}(x)|=9$         ($r=0.0875$)    $3.64\cdot 10^{-1}$    $3.41\cdot 10^{-1}$    $2.20\cdot 10^{-2}$    $3.48\cdot 10^{-1}$
#$\#|\text{B}_{r}(x)|=25$     ($r=0.13125$)    $4.51\cdot 10^{-5}$    $3.59\cdot 10^{-5}$    $3.95\cdot 10^{-5}$    $4.24\cdot 10^{-5}$
#$\#|\text{B}_{r}(x)|=49$          ($r=0.175$) $1.40\cdot 10^{-4}$ $\mathbf{1.39\cdot 10^{-5}}$ $2.93\cdot 10^{-5}$ $9.49\cdot 10^{-5}$
#$\#|\text{B}_{r}(x)|=81$     ($r=0.21875$)    $5.35\cdot 10^{-4}$    $1.01\cdot 10^{-4}$    $7.79\cdot 10^{-5}$    $1.76\cdot 10^{-4}$
#$\#|\text{B}_{r}(x)|=225$         ($r=0.35$) $3.92\cdot 10^{-3}$ $9.36\cdot 10^{-4}$ $4.28\cdot 10^{-3}$ $6.59\cdot 10^{-4}$

### v2

@@ Appendix E Finite Differences on Point Clouds
To find the stencil coefficients $c_{i}^{(x)}$ and $c_{i}^{(y)}$, we enforce that the approximation holds exactly for the functions $1$, $x$, and $y$, that is, we enforce that the approximation holds true for any polynomial of degree 1 in 2D. This results in the following systems of equations for the stencil coefficients $c_{i}^{(x)}$ and $c_{i}^{(y)}$,
$\sum_{i=1}^{N}c_{i}^{(x)}=0,\qquad\sum_{i=1}^{N}c_{i}^{(x)}(x_{i}-\tilde{x})=1,\qquad\sum_{i=1}^{N}c_{i}^{(x)}(y_{i}-\tilde{y})=0,$ (27)
@@
$\sum_{i=1}^{N}c_{i}^{(y)}=0,\qquad\sum_{i=1}^{N}c_{i}^{(y)}(x_{i}-\tilde{x})=0,\qquad\sum_{i=1}^{N}c_{i}^{(y)}(y_{i}-\tilde{y})=1.$ (28)
@@
$A\mathbf{c}^{(x)}=\mathbf{b}^{(x)},\qquad\qquad A\mathbf{c}^{(y)}=\mathbf{b}^{(y)},$ (29)
@@
$A=\begin{bmatrix}1&1&\cdots&1\\ x_{1}-\tilde{x}&x_{2}-\tilde{x}&\cdots&x_{N}-\tilde{x}\\ y_{1}-\tilde{y}&y_{2}-\tilde{y}&\cdots&y_{N}-\tilde{y}\end{bmatrix}$, (30)
@@
$\mathbf{c}^{(x)}=\begin{bmatrix}c_{1}^{(x)}\\ ...\\ c_{N}^{(x)}\end{bmatrix},\qquad\quad\mathbf{c}^{(y)}=\begin{bmatrix}c_{1}^{(y)}\\ ...\\ c_{N}^{(y)}\end{bmatrix},\qquad\quad\mathbf{b}^{(x)}=\begin{bmatrix}0\\ 1\\ 0\end{bmatrix},\qquad\quad\mathbf{b}^{(y)}=\begin{bmatrix}0\\ 0\\ 1\end{bmatrix}$. (31)
@@
$\mathbf{c}^{(x)}=(A^{\top}\!A)^{-1}A^{\top}\mathbf{b}^{(x)},\qquad\qquad\mathbf{c}^{(y)}=(A^{\top}\!A)^{-1}A^{\top}\mathbf{b}^{(y)}.$ (32)
Plugging these coefficients in Equation˜26 gives the desired approximation to the first-order derivatives at $f$.
[…]
@@
Appendix F Training Times of 2D mmGNOs with Autograd and Finite Differences
@@
We compare the training times per epoch when training 2D mmGNOs with $\mathcal{L}_{\mathrm{Poisson}}$ using Autograd, finite differences on uniform grids, and finite differences on non-uniform (NU) grids. Figures 16 and 17 display the training times per epoch versus the latent resolution and output resolution, respectively.
@@
As expected, we see from Figure˜16 that Autograd is more expensive than finite differences on a uniform grid at fixed latent and output resolutions. We can also see, by looking at individual columns corresponding to fixed latent space resolutions (i.e. the same model architecture), how finite differences at higher output resolutions compare to Autograd at lower output resolutions. In particular, non-uniform finite differences are often more expensive than Autograd at a slightly lower resolution. Figure˜17 quantifies how running time increases when the latent space resolution of the model increases.
@@
Figure 16: Time per epoch when training 2D mmGNOs with $\mathcal{L}_{\mathrm{Poisson}}$ using Autograd, finite differences on uniform grids, and finite differences on non-uniform (NU) grids. We display the training times per epoch at different latent space resolutions (x-axis) and different output resolutions (the numbers next to the points denote the number of points along each dimension).
@@
Figure 17: Time per epoch when training 2D mmGNOs with $\mathcal{L}_{\mathrm{Poisson}}$ using Autograd, finite differences on uniform grids, and finite differences on non-uniform (NU) grids. We display the training times per epoch at different output resolutions (x-axis) and different latent space resolutions (the numbers at the end of the curves denote the number of points along each dimension).
Appendix G Automatic Differentiation with Non-Differentiable Components
In our experiments, the entire mmGNO and mmGINO architectures are designed such that all of their components are differentiable, ensuring that derivatives of any order can be computed correctly and accurately using automatic differentiation. For instance, we avoid non-differentiable activation functions such as ReLU (Rectified Linear Unit), which introduce discontinuities in their derivatives, and instead use smooth alternatives like GeLU (Gaussian Error Linear Unit). All the other operations in the network, including the Fourier layers, pointwise linear neural networks, and pointwise linear operators, are also implemented in a differentiable manner to avoid issues related to non-differentiability.
While automatic differentiation is naturally suited to smooth functions, it can accommodate non-differentiable components. For example, ReLU is non-differentiable at 0, yet Autograd computes exact gradients on each side and assigns a subgradient at 0 (e.g. 0 or 1). This works well in practice since exact zero inputs are rare and subgradient-based optimization remains effective. Other strategies exist for non-smooth components. Clarke subderivatives formalize generalized derivatives at non-smooth points. Proximal methods split updates into a smooth gradient step and a proximal mapping for the non-smooth part. Randomized smoothing adds small perturbations to inputs and averages outputs. Straight Through Estimators use the true non-differentiable function in the forward pass while replacing its gradient in the backward pass with a differentiable surrogate, such as a smoothed function or the identity.
Physics-informed approaches use automatic differentiation to compute derivatives of network outputs with respect to input coordinates. While automatic differentiation can handle non-differentiable components, they can be problematic for physics-informed losses. Non-smoothness can create regions where derivatives are undefined or discontinuous, and these issues can propagate through the chain rule when higher-order derivatives are needed for PDE residuals. The result can be abrupt residual changes, numerical instability, and convergence to suboptimal solutions. Differentiable components are therefore preferable since they provide continuous derivatives, ensuring smooth residuals that better capture the physics. This reduces numerical artifacts, improves gradient quality, and supports stable optimization, leading to faster convergence and more accurate solutions, with greater reliability.